A new conveyor system at the local packaging plan will utilize a motor-powered mechanical arm to exert an average force of 890 N to push large crates a distance of 12 meters in 22 seconds. Determine the power output required of such a motor.


1
Expert's answer
2021-02-23T10:02:28-0500

The power of a motor

"P=\frac{work\: done}{time}=\frac{W}{t}=\frac{Fd}{t}"

"P=\rm\frac{890\: N\times12\: m}{22\: s}=485\: W"